HUBBLE SPACE TELESCOPE DAILY REPORT #2624 PERIOD COVERED: 0000Z (UTC) 05/15/00 - 0000Z (UTC) 05/16/00 Daily Status Report as of 137/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.2 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.3 Completed STIS/CCD/MA1/MA2 8424 (MAMA Sensitivity and Focus Monitor C8) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of object BD+28D4211 in order to measure the sensitivity and to focus monitor C8. This also is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes, and to monitor the STIS focus in a spectroscopic mode. A wavelength calibration was also taken. The observations were completed as planned, and no anomalies were reported. 1.4 Completed Three Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working. The proposal completed with no reported problems. 1.5 Completed Six Sets of WF/PC-2 8442 (Cycle 8 Standard Darks) The WF/PC-2 was used to obtain dark calibration dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported anomalies. 1.6 Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.7 Completed WF/PC-2 8459 (Decontaminations and Associated Observations Pt. 2/2) The WF/PC-2 was used to perform the monthly decons. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, and UV throughput checks. The observations were completed nominally. 1.8 Completed Three Sets of STIS/CCD 8808 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted, parallel observations as part of a POMS test proposal. The observations completed with no anomalous activity. 1.9 Completed Two Sets of STIS/CCD 8798 (Read Noise Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to measure the read noise of the STIS CCD using pairs of bias frames. The proposal completed nominally. 1.10 Completed FGS 8292 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) The Fine Guidance Sensor (FGS1R) was used to make astrometric observations of the star GJ1215 and other nearby reference stars. This is part of a study to calibrate the mass-luminosity relation for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. The observations were completed as planned, and no problems were reported. 1.11 Completed Two Sets of WF/PC-2 8346 (Cosmological Parameters from Type Ia Supernovae at High Redshift) The WF/PC-2 was used to directly measure the cosmological parameters, Omega, Lambda, and thus the curvature, Omega_k, using Type Ia supernovae {SNe Ia} as calibrated standard candles. Observations of SNe Ia spanning a large redshift range are key to decoupling Omega and Lambda. The observations completed nominally. 1.12 Completed Three Sets of STIS/CCD 8796 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV) The Space Telescope Imaging Spectrograph (CCD) was used in the parallel mode to make some low galactic latitude archive observations. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.14 Completed FOC/96 7923 (FOC F/96 Turn-on) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed STIS/CCD 8489 (Searching for the Hydrogen Reionization Edge of the Universe at 5HST) The Space Telescope Imaging Spectrograph (CCD) was used to observe 36 parallel orbits {4--5 fields of 5--8 orbits each} to constrain the H Lyman-edge in emission that marks the transition from a neutral to a fully ionized IGM. This edge is due to recombination from the H Lyman series and Lyman continuum, and can be used to constrain zion, one of the most important unknown quantities in large scale structure and cosmology. The proposal completed nominally. 1.16 Completed STIS/CCD/MA1 8097 (C IV High Velocity Clouds: Remnants of the Local Group?) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of two high clouds {HVCs} in the direction of Mrk 509 that have unusual ionization properties. The clouds exhibit strong C IV absorption with little or no detectable low ion {C II, Si II} absorption or H I 21cm emission. The observations completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 9 Successful: 9 2.2 FHST Updates: Scheduled: 12 Successful: 12 2.3 Operations Notes: The SSR EDAC error counter was reset once per ROP SR-1A. Using ROP IC-1, the MA transponder was commanded twice. The CCS "A" string pipeline was re-cycled at 136/1342Z. The ephemeris tables were uplinked at 137/0241Z as directed by ROP DF-07B. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW